OpenStudy (anonymous):

Solve for i 9x-7i > 3(3x-7u)

OpenStudy (igreen):

First distribute 3 into the parenthesis. 3 * 3x = ? 3 * -7y = ?

OpenStudy (igreen):

-7u*

OpenStudy (igreen):

Do you know how to multiply those?

OpenStudy (anonymous):

no

OpenStudy (igreen):

Just multiply the numbers together and keep the variable. For example: 9 * 7x = 9 * 7 = 63x

OpenStudy (igreen):

Get it? So can you do 3 * 3x?

OpenStudy (anonymous):

is it 9x?

OpenStudy (anonymous):

sorry i suck at math

OpenStudy (igreen):

Yes, you got it. How about 3 * -7u? @EmmanuelRocksYou

OpenStudy (anonymous):

-21u?

OpenStudy (igreen):

Yep

OpenStudy (igreen):

So we have: 9x - 7i > 9x - 21u Subtract 9x to both sides, what's 9x - 9x?